OpenStudy (anonymous):

Which equation is a quadratic equation? y = −4(3x + 2) y + 3x3 = (x − 2)(3x + 8) y − 3 = (2x2 + 11)(x − 1) + 3x y + 5x = −2x(4 − x) + 1

jimthompson5910 (jim_thompson5910):

Which one do you think it is?

jimthompson5910 (jim_thompson5910):

y = −4(3x + 2) y = −12x - 8 which is in y = mx+b form, which is a line, so that rules choice A out

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

you want something where you can rearrange it into the form y = ax^2+bx+c

jimthompson5910 (jim_thompson5910):

so which answer choice can be written into y = ax^2+bx+c form

OpenStudy (anonymous):

so either C or D right?

jimthompson5910 (jim_thompson5910):

good, choice B is ruled out because it has an x^3 term

jimthompson5910 (jim_thompson5910):

what happens when you multiply out (2x^2 + 11)(x − 1)

OpenStudy (anonymous):

i don't know

jimthompson5910 (jim_thompson5910):

what is 2x^2 times x

OpenStudy (anonymous):

2x^3

OpenStudy (amoodarya):

look at this image

jimthompson5910 (jim_thompson5910):

so that rules out choice C because choice C has a cubic term

OpenStudy (anonymous):

oh thanks for helping me

jimthompson5910 (jim_thompson5910):

np

OpenStudy (amoodarya):

so D is right